### v2

# Parameter inference of microlensed gravitational waves using neural spline flows

Abstract
When gravitational waves (GWs) propagate near massive objects, they undergo gravitational lensing that imprints lens model dependent modulations on the waveform. This effect provides a powerful tool for cosmological and astrophysical studies. Due to the added parameters of lenses and the uncertainty of lens models, parameter inference for lensed GW events using traditional methods is extremely time-consuming, thus requiring more efficient parameter inference methods. In this work, we explore the use of neural spline flows (NSFs) for posterior inference of microlensed GWs, and successfully apply NSFs to the inference of 11-dimensional lens parameters. Our results demonstrate that compared with traditional methods like Bilby dynesty that rely on Bayesian inference, the NSF network we built not only achieves inference accuracy comparablye to traditional methods for the main parameters ($M_{\rm L}$, $y$, $\mathcal{M}_{\rm c}$, $d_{\rm L}$), but also can reduce the inference time from approximately 3 days to 0.8 s on average. Additionally, the network exhibits strong generalization for the spin parameters of GW sources. It is anticipated to become a powerful tool for future low-latency searches for lensed GW signals.
I Introduction
Since the first direct detection of gravitational waves (GWs) in 2015, LIGO-Virgo-KAGRA (LVK) network have reported $93$ GW events in its first three observing runs, which enables us to conduct various novel tests of general relativity Abbott et al. (2016, 2018); Abbott et al. (2019a, b, 2021a) and cosmological researches The LIGO Scientific Collaboration et al. (2021); Abbott et al. (2023). With the further analysis of the data observed in the fourth observing run of the LVK network and the construction of various high-performance gravitational wave detectors in the future (such as the Laser Interferometer Space Antenna (LISA) Barausse et al. (2020), Taiji Ruan et al. (2020), TianQin Luo et al. (2025), Einstein Telescope (ET) Maggiore et al. (2020), and Cosmic Explorer (CE) Evans et al. (2021)), the number of detected GW events will increase sharply, providing a brand-new powerful tool for research in fields such as fundamental physics, cosmology, and astrophysics. Among these GW signals, it is particularly worthy of attention those signals affected by the lensing effect. The lensing effect of GWs is caused by celestial bodies located between detectors and sources, which will result in a significant modulation of GW signals Dai and Venumadhav (2017); Dai et al. (2018).
As an important tool in GW astronomy, the observation of lensed GW signals holds great research value. Some celestial bodies (such as dark matter halos Choi et al. (2021); Guo and Lu (2022), intermediate-mass black holes Lai et al. (2018); Gais et al. (2022); Meena (2024) and cosmic strings Fernández-Núñez and Bulashenko (2016, 2017)) neither emit light nor neutrinos, thus cannot be detected by conventional observational means. However, if GWs are lensed by them, the detection of such lensed signals may offer an effective method for the observation and study of these celestial bodies, and even lead to the discovery of new ones. Particularly important is that the detection of lensed GW signals is expected to provide direct probe for the nature of dark matter Liu et al. (2019); Jung and Shin (2019); Urrutia and Vaskonen (2022); Wang et al. (2021a); Guo and Lu (2022); Cao et al. (2022); Fairbairn et al. (2023); Çalışkan et al. (2023); Tambalo et al. (2023a); Cheung et al. (2024); Jana et al. (2024). Not only those, this effect is also expected to provide us with some more precise tests in cosmology Cao et al. (2019); Liu et al. (2022); Goyal et al. (2023); Huang et al. (2023); Smith et al. (2025).
Recently, multiple studies have conducted numerous searches for the GW lensing effect, but none have clearly detected lensed GW signals. The methods employed include traditional search techniques based on matched filtering Abbott et al. (2021b); Abbott et al. (2024); Janquart et al. (2023); Barsode et al. (2025); Goyal et al. (2024) and those incorporating deep learning technology Goyal et al. (2021); Kim et al. (2022); Schäfer et al. (2023). Nevertheless, as more and more GW events are expected to be reported, and given the predictions regarding the detection rate Li et al. (2018); Ng et al. (2018); Wang et al. (2021b); Xu et al. (2022); Brando et al. (2025), we remain hopeful that GW signals affected by the gravitational lensing effect will be detectable in future observational runs.
Due to the difference in wavelength, the gravitational lensing effect of GWs is distinct from that of electromagnetic waves (EMWs). The wavelength of EMWs is usually much smaller than the Schwarzschild radius of the lensing object, and the geometric optics (GO) approximation can generally be directly applied. The observation frequency range of ground-based GW detectors is very low, compared with the typical frequencies of EMWs. For GW signals detected by the LVK network, their wavelengths are usually larger than the Schwarzschild radius corresponding to lensing objects with mass $M_{\rm L}\lesssim 10^{3}$ ${\rm M}_{\odot}$, leading to non-negligible wave effects Meena and Bagla (2020); Ali et al. (2023). Therefore, when studying lensed GWs, it is necessary to fully consider whether the GO approximation is valid Takahashi and Nakamura (2003).
When GWs are affected by lenses of different masses, they will undergo different waveform changes. When the mass of the lens is extremely large, repetitive GW events will occur at intervals of several minutes or even months, forming multiple “images” of GW sources, which is the strong gravitational lensing effect Sereno et al. (2010). In contrast, the microlensing effect caused by stars or compact objects has repetitive event intervals of only a few milliseconds to a few seconds. Under the microlensing effect, the time delay caused by the gravitational potential of the lens during the propagation of GWs is comparable to the period of the wave, which not only amplifies the signal but also generates a “beating” pattern Bulashenko and Ubach (2022) in the waveform. By using microlensed GW signals as probes, we will be able to conduct more comprehensive studies on compact objects such as stars and small-mass black holes that act as lenses.
Parameter inference for lensed GW events is a crucial step in studying the GW lensing effect. The traditional parameter inference process based on Bayesian methods is extremely time-consuming, especially for lensed GW events with additional lens parameters, which significantly reduces its efficiency. Recent studies have shown that deep learning techniques such as Conditional Variational Autoencoders (CVAE) Gabbard et al. (2022) and Normalizing Flows (NFs) Green and Gair (2021); Green et al. (2020); Dax et al. (2021); Shen et al. (2022); Whittaker et al. (2022); Dax et al. (2023); Zhao et al. (2023); Sun et al. (2024a); Du et al. (2024); Sun et al. (2024b); Leyde et al. (2024); Dax et al. (2025); Xiong et al. (2025) can effectively and rapidly perform parameter inference, providing powerful tools for low-latency GW signal searches. In Ref. Nerin et al. (2025), the use of CVAE for parameter inference of microlensed GW events was investigated, achieving rapid parameter inference in two parameters, source’s dimensionless location parameter $y$ and time delay $\Delta t_{\rm d}$, with the fastest inference speed being $5$ orders of magnitude faster than the Bayesian methods such as Bilby. In this work, we focus on microlensing events, using a point-mass lens model to describe the lensing effect on GWs, and explore the use of neural spline flows (NSFs) for parameter inference in 11 dimensions (including both sources and lenses information, i.e., lens mass $M_{\rm L}$, angular diameter distance of lens $D_{\rm L}$, source chirp mass $\mathcal{M}_{\rm c}$, mass ratio $q$, luminosity distance of source $d_{\rm L}$, location parameter of the source $y$, merger phase $\phi_{\rm c}$, merger time $t_{\rm c}$, orbital inclination $\iota$, right ascension $ra$, and declination $dec$).
GW sources (such as the merger of binary black holes) typically exhibit spin effects Ajith et al. (2011); Callister et al. (2022); Biscoveanu et al. (2022); Fishbach et al. (2022), and studying their spins is of great significance Roulet et al. (2021); Biscoveanu et al. (2021); Tong et al. (2022). However, traditional Bayesian methods encounter significant challenges when dealing with waveforms containing spin parameters Veitch et al. (2015); Harry et al. (2016); Shi et al. (2025). Utilizing the rapidity Du et al. (2024); Wang et al. (2024a) and generalization ability Chatterjee et al. (2021); Xia et al. (2021); Wang et al. (2024b); Xu et al. (2024); Wang et al. (2024c) of the NSF network can also provide a solution to this difficulty.
In this work, we trained an NSF network on a set of simulated microlensed GW datasets, enabling it to learn the probabilistic latent representation of the lensing system parameters based on the GW time series signals. The results show that the network can effectively perform posterior inference for microlensed GW events. Additionally, we demonstrate that the network has good generalization for spin parameters not introduced during training, further expanding the applicability of the network.
The organization of this paper is as follows. Section  II provides a detailed introduction to the methodology used in our work, including a review of the gravitational lensing theory, the configuration of the network, and the data generation part. Section  III presents the obtained results. Finally, Section  IV summarizes the conclusions of this work and proposes improvement directions and the next research plan.
II METHODOLOGY
II.1 The theory of gravitational lensing
In the gravitational lensing effect, due to the impact of a celestial body known as lens on GWs, we will observe the modulation of signals. As shown in Fig. 1, this effect occurs when the source is located near the axis determined by the centers of lens and observer. Since the size range of lens and source is much smaller than the distances from source to observer and from lens to source, they can be projected onto the lens plane and the source plane respectively Schneider et al. (1992) for analysis. $D_{\rm L}$, $D_{\rm LS}$, and $D_{\rm S}$ are angular diameter distances from observer to lens, from lens to the source, and from the source to the observer, respectively. The vector $\bm{\eta}$ represents the location of the GW source relative to the axis on the source plane, and the vector $\bm{\xi}$ represents the impact parameter of lens on the GW propagating to the lens plane. Under this thin-lens approximation, GWs are only affected by two-dimensional gravitational potential $\psi(x)$ determined by the properties of lens on the lens plane. This impact is ultimately reflected in the transmission factor $F(f)$ dependent on the lens model. Here, $F(f)=\tilde{h}^{L}(f)/\tilde{h}(f)$ connects the lensed waveform $\tilde{h}^{L}(f)$ and the unlensed waveform $\tilde{h}(f)$. Therefore, in waveform data simulation stage, we only need to calculate $\tilde{h}(f)$ and $F(f)$ separately first, and then multiply the two to generate the waveform data of the lensed GWs.
Figure 1: A classical gravitational lensing system composed of a GWs source on the source plane, a lens on the lens plane, and the observer. $D_{\rm L}$, $D_{\rm LS}$, $D_{\rm S}$ are angular diameter distances from observer to lens, from lens to source, and from source to observer, respectively. The vector $\bm{\eta}$ represents the GWs source’s location. The vector $\bm{\xi}$ is the impact parameter of the lens on GWs.
For axisymmetric lenses, the vectors $\bm{\eta}$ and $\bm{\xi}$ can be simplified to scalars $\eta$ and $\xi$. On the source plane, a dimensionless scalar $y\equiv\eta/(\xi_{0}D_{\rm S}/D_{\rm L})$ is taken as the location parameter of the source, where $\xi_{0}$ is a characteristic length on the lens plane that depends on the lens model. In the point mass lens (PML) model, it is chosen as the Einstein radius:
$\xi_{0}=\sqrt{\frac{4GM_{\rm L}}{c^{2}}\cdot\frac{D_{\rm L}D_{\rm{LS}}}{D_{\rm S}}}.$ (1)
The PML model approximates the lens as a point mass and is applicable when the physical size of the lens is much smaller than the Einstein radius, such as black holes, dense dark matter clumps, and similar compact celestial bodies.
The gravitational lensing effect will cause repeated GW events with certain time delay at the observer’s location. For the PML model, under the GO approximation condition, two “images” of the GW sources will be produced at the observer’s location Ali et al. (2023). At this point, $F(f)$ is given in Ref. Takahashi and Nakamura (2003) as:
$F(f)=|\mu_{+}|^{1/2}-i|\mu_{-}|^{1/2}e^{2\pi if\Delta t_{\rm d}},$ (2)
where
$\mu_{\pm}=\frac{1}{2}\pm\frac{y^{2}+2}{2y(y^{2}+4)^{1/2}},$ (3)
$\Delta t_{\rm d}=\frac{2GM_{\rm Lz}}{c^{3}}\left\{y(y^{2}+4)^{1/2}+2\ln\left[\frac{(y^{2}+4)^{1/2}+y}{(y^{2}+4)^{1/2}-y}\right]\right\},$ (4)
are the magnification factors and time delay for the two “images” respectively. $M_{\rm Lz}=M_{\rm L}(1+Z_{\rm L})$ is redshift mass of the lens. Thus, the impact of the gravitational lensing effect can be determined by $y$ and $M_{\rm L}$, and further reflected by transmission factor $F(f)$.
II.2 Network architecture
This work proposes a parameter inference network based on deep residual feature extraction and NSFs for efficiently reconstructing the 11-dimensional posterior distribution of physical parameters from simulated multi-detector lensed GW signals. The network is composed of a feature encoder and a probabilistic decoder cascaded together.
The feature encoder adopts an improved ResNet-50, whose structure is exactly the same as that described in Ref. He et al. (2016). Its input is the four-channel time-domain signal, corresponding to the GW signals detected by the four detectors: LIGO (H1/L1), Virgo (V1), and KAGRA (K1). The time-domain features are ultimately mapped into a 512-dimensional feature vector through the ResNet-50 network, and then it is transmitted as input to the probabilistic decoder for posterior inference.
The probability decoder adopts the NSF architecture, specifically the Rational Quadratic Neural Spline Flows (RQ-NSFs) Durkan et al. (2019). The NF is a probabilistic modeling method that utilizes deep learning techniques to learn complex probability distributions. Its core idea is to construct a bijective transformation $f_{x}$ from parameters to dependent variables, thereby obtaining the mapping relationship between a simple base distribution (such as a normal distribution) and a complex posterior distribution. It is a reversible neural network capable of accurately calculating the Jacobian determinant of the probability density function through network transformations. Each transformation is based on the input data and the output of the previous layer of the neural network, thereby predicting and gradually approximating the posterior distribution. RQ-NSF enhances the flexibility of autoregressive transformations by introducing monotonic rational quadratic splines in NFs. It uses monotonic rational quadratic transformations to replace the addition or affine transformations in the regression layer and reverses the sample components between autoregressive transformations to ensure that the transformation sequence can specifically and completely transform all components. Through the invertible and differentiable transformation $f_{x}$, the parameter posterior distribution derived by the NSF network can be analytically expressed as Rezende and Mohamed (2015); Papamakarios et al. (2017):
$p(\theta|x)=\pi\left(f_{x}^{-1}(\theta)\right)\left|\det J_{f_{x}}^{-1}\right|,$ (5)
where $x$ represents the data input to the network, and $\theta$ corresponds to the parameters to be inferred. The $u=f_{x}^{-1}(\theta)$ is linked to $\theta$ through the transformation $f_{x}$, and $\pi(u)$ is usually chosen to be a standard normal distribution.
Network training is conducted using the maximum likelihood criterion, with the optimization objective being to minimize the negative log-likelihood loss:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\log p(\theta_{i}|x_{i}).$ (6)
The NSF network’s key parameters are shown in Table 1. During the training process, to ensure the validity of the predictions made by the NSF network and prevent overfitting, we employ the AdamW optimizer and adopt a 50-round early stopping strategy. We also regenerate $1.2\times 10^{4}$ new data points at the beginning of each training cycle, with $10^{4}$ used for training and $2\times 10^{3}$ for validation. The training process of the network is illustrated in Fig. 2. Our entire network is trained on an NVIDIA RTX A6000 GPU with 48 GB of memory.
Table 1: The key parameters of the NSF network.
Parameter Value
Activation function ReLU
Number of bins 8
Flow steps 3
Transform blocks 5
Batch size 20
Learning rate decay factor $2\times 10^{-5}$
Initial learning rate $2\times 10^{-4}$
Hidden layer size 1024
Figure 2: Workflow of the NSF network. The input is the GW time series signal after lensing. The signal passes through the ResNet-50 network (that is the part enclosed by the blue dotted line) to extract feature vectors. Finally, the extracted feature vectors are inputted into the NSF network (that is the part enclosed by the green dotted line) to obtain posterior inference of parameters.
II.3 Data sets
The training, validation and test sets of the network are both composed of parameter sets and the corresponding GWs time-domain signals. The data are based on a flat $\Lambda$CDM cosmological model ($H_{0}=70$ km/s/Mpc, $\Omega_{m}=0.3$, $\Omega_{\Lambda}=0.7$), and are generated using appropriate prior parameters (see Table 2).
Table 2: The prior ranges of each parameter used to generate GWs waveform data. For the convenience of the research, the parameters not mentioned in the table have been simply set to zero. Only when studying the generalization ability of the network for spin parameters, we change the spin values to generate a series of test sets containing spins.
Parameter Description Range Unit Distribution
$\mathcal{M}_{\rm c}$ Chirp mass of the binary system $(5,80)$ $\mathrm{M}_{\odot}$ LogUniform
$q$ Ratio of masses of the two black holes $(0.125,1)$ Uniform
$M_{\rm L}$ Mass of the lens $(10^{3},10^{5})$ $\mathrm{M}_{\odot}$ LogUniform
$Z_{\rm L}$ Redshift value at the lens $(0,1)$ PowerLaw
$Z_{\rm S}$ Redshift value at the source $(Z_{\rm L},1)$ PowerLaw
$\delta$ Ecliptic latitude of the binary system $\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$ $\mathrm{rad}$ Cosine
$\alpha$ Ecliptic longitude of the binary system $(0,2\pi)$ $\mathrm{rad}$ Uniform
$\psi_{\rm s}$ Polarization angle of the gravitational wave $(0,\pi)$ $\mathrm{rad}$ Uniform
$\iota$ Angle of inclination of the binary orbit $(0,\pi)$ $\mathrm{rad}$ Sine
$\phi_{\rm c}$ Phase at the moment of coalescence $(0,2\pi)$ $\mathrm{rad}$ Uniform
$t_{\rm c}$ Time of coalescence $(7,8)$ $\mathrm{s}$ Uniform
$\eta$ Location of the binary system $(10^{-6},0.5)$ $\mathrm{pc}$ Uniform
In addition, the prior of $d_{\rm L}$ from the source to observer and $D_{\rm L}$ can be obtained from $Z_{\rm L}$, and both follow a power-law distribution. The probability density function for $Z_{\rm L}$ and $Z_{\rm S}$ are as follows: $p(Z_{\rm L})\propto 3Z_{\rm L}^{2}$, and $p(Z_{\rm S}|Z_{\rm L})\propto 3/(1-Z_{\rm L}^{3})Z_{\rm S}^{2}$. To better cover the parameter space and achieve uniform sampling at different orders of magnitude, the selected $M_{\rm L}$ and $\mathcal{M}_{\rm c}$ follow a log-uniform distribution. The selection of $M_{\rm L}$ and $\eta$ range mainly refers to Ref. Kim et al. (2022). It is indicated in Ref. Takahashi and Nakamura (2003) that when $M_{\rm Lz}$ and the frequency $f$ of GW satisfy the condition that $\omega\equiv 8\pi GM_{\rm Lz}f/c^{3}\gtrsim 1$, transmission factor $F(f)$ in the wave optics (WO) asymptotically converges to the GO limit. For the cases $M_{\rm L}>10^{3}$ ${\rm M}_{\odot}$ and $f>20$ ${\rm Hz}$ considered in our study (correspondingly $\omega>7.5$), GO approximation is applicable within a certain range. Although the most rigorous approach would be to directly adopt the analytical form of $F(f)$ Takahashi and Nakamura (2003), for the purpose of exploratory verification and considering the computational cost and practicality, we chose a more concise GO approximation result. For discussions on the applicable range of the geometric optical approximation in the lensing effect of GWs, refer to Refs. Meena and Bagla (2020); Bulashenko and Ubach (2022); Bondarescu et al. (2023); Takahashi (2004).
[…]
The training set, validation set and test set are generated using the same method, and these datasets are mutually exclusive. Both the normal training set and the test set are generated without spin. To test the network’s generalization ability for spin parameters, we also generate a series of other test sets with different spin values using the same method. Before inputting the parameter set samples into the network for training, we normalize them to the range of $[-1,1]$ using the max-min normalization method.
III Results and discussion
We train the NSF network using the architecture described in Section  II.2. After $257$ epochs, the loss function achieves a stable evolution, as shown in Fig. 3. The loss value of training set eventually converges to about $-4.7$, and the training loss and validation loss remain close throughout, with no sign of overfitting. At the end of training, we save the network of the epoch with the best performance, that is, the one with the lowest validation loss, and use it for testing.
We randomly select $200$ events from the test set to verify the robustness of the NSF network’s inference for selected 11 dimensions parameters. By injecting $200$ simulated lensed GW waveforms, we measure the frequency of the true parameters at a series of confidence levels. We also calculate the cumulative probability of the percentile of each parameter’s posterior distribution at the injected parameter value, and then plot the cumulative distribution function (CDF) of the joint distribution of the parameters in Fig. 4. Ideally, the injected parameter CDF should show a diagonal pattern because the CDF represents the cumulative sum of the distribution of each event. The closeness of the curve to the diagonal indicates the accuracy of the parameter inference. As shown in the figure, the curve closely follows the diagonal, and the entire curve is within the $95\%$ confidence interval, indicating that the NSF network perform excellently in accurately estimating parameters. The $p$-value obtained from the posterior distribution of the network is also marked in the legend.
Figure 3: Changes in loss values on the training set and validation set during training. The NSF network is finally converged quite well.
Figure 4: Cumulative distribution of the quantiles where the true values lie in the edge distribution of the NSF network. This curve is obtained using 200 samples, and the dotted black diagonal line represents the ideal result. The legend indicates the corresponding $p$-value, and the gray area represents the $95\%$ confidence interval. It demonstrates that the parameter inference accuracy of the NSF network is quite satisfactory.
Figure 5: Marginal one-dimensional and two-dimensional posterior distribution plots for the same event without spin, inferred using the NSF network and Bilby dynesty. The true values of the parameters are marked by black dashed lines. The blue curves represent the results from the NSF network, while the red curves represent the results from Bilby dynesty. In the two-dimensional distribution plots, the contour lines represent the $95.4\%$ and $68.3\%$ confidence regions, respectively. It demonstrates the accuracy of NSFs in restoring the key parameters, as well as its ability to identify the degeneracy of some parameters.
The $p$-value provides a quantitative measure of the consistency between the cumulative distribution predicted by the network and the expected theoretical distribution. It is calculated based on the Kolmogorov-Smirnov (KS) test Lopes (2011) method. In GW parameter estimation, the KS test can be used to assess the degree of agreement between the estimated parameter distribution and the expected theoretical distribution Biwer et al. (2019). The KS test first calculates the CDF of the two distributions and then computes the distance which is called the KS statistic between them based on the CDF. The $p$-value can be calculated based on the KS statistic and is typically used to quantitatively evaluate the similarity or difference between two distributions. By comparing the $p$-value with the significance level, if the $p$-value is greater than the significance level, it indicates that the two compared distributions are statistically similar. Here, the significance level is set at 0.05.
We evaluate the reasonableness of the posterior established by the network by comparing two distributions, namely the ideal distribution and the joint distribution of parameters generated by the NSF network. The ideal distribution is a uniform distribution because the percentiles are uniformly distributed between 00 and $1$. The other distribution is the cumulative probability of the percentiles of each parameter’s posterior distribution at the parameter injection value. Our null hypothesis is: assuming that these two distributions are the same, this means that the posterior inference of the parameters given by the NSF network is reasonable; while the alternative hypothesis naturally states that the two distributions are different, meaning that the posterior given by the network is unreasonable. The calculated $p$-value is compared with the significance level of $0.05$. If the $p$-value is less than $0.05$, the null hypothesis is considered invalid. the NSF network we trained ultimately gives a $p$-value $=0.5517>0.05$, which indicates that the network’s inference of the posterior is reasonable.
Figure 5 shows a comparison of the posterior inference of a lensed GW event in the spinless test set sample between our trained NSF network and Bilby dynesty, where the prior settings of Bilby dynesty are the same as those in Table  1. For the 11 parameters we studied, the NSF network successfully completed the posterior inference in less than $1$ s and provided relatively accurate posterior estimates for the lens parameters $M_{\rm L}$ and $y$, which are of the greatest concern to us. Here, $M_{\rm L}$ refers to the lens mass in the source frame (not the red-shifted mass). Meanwhile, it also gave good inference results for some parameters of the source, such as $\mathcal{M}_{\rm c}$ and $d_{\rm L}$. For parameters $\iota$, $q$, $\delta$ and $\alpha$, NSFs provided a bimodal distribution, which was not obvious when compared with the results of Bilby dynesty in the figure. NSFs performed poorly for $\psi_{\rm s}$, $\phi_{\rm c}$ and $t_{\rm c}$, and we believe that this is due to the further modulation of GWs by the lensing effect, which increases the difficulty of extracting features related to the source, preventing NSF network from extracting these parameter features well. Overall, the two-dimensional posterior distribution given by NSFs is relatively diffuse. We believe this is because the network did not extract the high-order mode features Estellés et al. (2022a, b); Mateu-Lucena et al. (2022); Diao et al. (2025) introduced by the IMRPhenomTPHM model well.
In contrast, the inference accuracy of Bilby dynesty for the parameters $M_{\rm L}$, $y$, $\mathcal{M_{\rm c}}$ and ${d}_{\rm L}$ is not better than that of the NSF network. Due to the introduction of higher-order modes of the waveform, the two-dimensional posterior distribution given by Bilby dynesty is indeed more concentrated. Although it had relatively better accuracy in inferring other parameters, the time it took is four orders of magnitude longer than that of NSFs. In our $10$ tests (see Table 3), the posterior inference speed of Bilby dynesty is the fastest at $60070$ s, while the inference speed of the NSF network averaged approximately $0.8$ s. This makes it possible for NSFs to apply in real-time GW signals search.
Table 3: The statistical results of the time taken for Bilby dynesty and NSFs network to independently complete the posterior inference of 11-dimensional parameters, respectively, after $10$ tests conducted.
NSFs Bilby dynesty
(s) (s)
Average 0.8 256490
Maximum 0.8 1471691
Minimum 0.8 60070
Figure 6 shows the posterior inference performance of the NSF network on the test set that includes different spin parameters. This figure depicts the distribution of error bars for $p$-values under $36$ spin values. For each spin value, we randomly select one sample from the corresponding test set (with a sample size of 2000) and perform $16$ random samplings (each sampling containing 200 events) to calculate $16$ $p$-values. These original $p$-values are represented by light gray dots in the figure. To more clearly show the trend of network performance with respect to spin values and to eliminate the influence of injected noise, we divide $p$-values under $36$ spin values into groups of $4$ for binning statistics and draw error bars. The center position of the error bars represents the mean of $p$-values in this group, and their length indicates the range of the $95\%$ confidence interval. The results show that when the spin value is less than approximately $0.17$, $p$-values are generally higher than $0.05$, indicating that the NSF network has good generalization ability within this range. However, as the spin value increases, $p$-values show an overall downward trend. When spin value is greater than $0.17$, $p$-values are significantly lower than $0.05$, which can be regarded as the generalization boundary of the NSF network for the assumed spin value scenario under the current training conditions.
Figure 6: The KS test evaluating the performance of the NSF network on the test set with added spins. We assume equal spins for the two black holes ($\chi_{\rm 1z}=\chi_{\rm 2z}$), which increase incrementally from 00 to $0.35$. The dot at the center of each error bar represents the mean $p$-value after uniformly binning the 36 datasets. The blue line illustrates the trend of these $p$-values across spin values. The light gray dots in the background represent $p$-values separately calculated for the 36 spin values, and the color shade indicates the degree of overlap of $p$-values. The black dotted line indicates a significance level corresponding to a $p$-value of $0.05$.
IV Conclusion and outlook
In this work, we investigated the applicability of the NSF network in parameter inference for microlensed GW events. Specifically, we trained an NSF network to infer the 11-dimensional parameters describing the lens and source in the point-mass lens model. For a randomly selected event, the network infers parameters $M_{\rm L}$, $D_{\rm L}$, $y$, $\mathcal{M}_{\rm c}$, $D_{\rm L}$, $q$, $\delta$, $\alpha$ and $\iota$ not only much faster than Bilby dynesty, but also did not differ too much in accuracy. Although the accuracy of inferring parameters $\psi_{\rm s}$, $\phi_{\rm c}$, and $t_{\rm c}$ is not ideal, it can still rapidly narrow down the estimated range of these parameters. Moreover, the NSF network demonstrates good generalization for spin parameters not included in the training, indicating that it may be capable of adapting to a wider variety of events.
In conclusion, this study demonstrates that NSFs can significantly enhance the search efficiency within the parameter space and can efficiently complete the parameter inference for microlensed GW events.Compared with the traditional Bayesian method, NSFs’ average inference speed in a 12-dimensional parameter space is four orders of magnitude faster than that of Bilby dynesty. This makes it promising to provide strong support for low-latency searches of lensed GW events.
It is worth noting that although we just adopt PML model in this paper, there exist many uncertainties in lens models. Considering the uncertainties on lens models, much more extra parameter space needs to be searched. The NSF network provides a much more rapid method for parameter inference under a very high dimension parameter space including lens models.
What’s more, for the purpose of simplifying the calculation, our current research adopts a simple GO approximation to calculate $F(f)$. We can expand it to the range of WO or adopt other more precise calculation methods Guo and Lu (2020); Tambalo et al. (2023b) in the future. For example, in Ref. Nerin et al. (2025), by using WO at low frequencies and GO approximation at high frequencies, a certain degree of balance between calculation efficiency and accuracy is achieved. However, in any case, the simulation process for lensed waveform data is similar.
In the future, we will further investigate why the NSF network is not accurate enough in inferring parameters such as $\psi_{\rm s}$, $\phi_{\rm c}$, and $t_{\rm c}$. This may involve improving the architecture of the NSF network itself. We will also explore the use of the NSF network’s ability to quickly narrow down the inference range of these parameters, and use the network’s posterior estimates of these parameters as the prior for bilby to complete the entire parameter inference process Williams et al. (2021); Nerin et al. (2025). This will help us explore the combined use of the NSF network and traditional parameter inference methods. Additionally, for most lensed GW signals, both strong lensing and microlensing effects are present simultaneously. Some investigations have shown that strong lenses make microlensing effects easier to detect Diego (2020); Mishra et al. (2021); Cheung et al. (2021); Seo et al. (2022), and further research on the parameter inference effect under this combined effect will be more practical and valuable.
DATA AVAILABILITY
The data that support the findings of this article are not publicly available upon publication because it is not technically feasible and/or the cost of preparing, depositing,and hosting the data would be prohibitive within the terms of this research project. The data are available from the authors upon reasonable request.
SOFTWARE AVAILABILITY
The code used for simulating GW and data processing techniques is used by Pycbc. The code for lensing the simulated GW can be requested from the corresponding authors upon reasonable request.
Acknowledgements.
This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 12473001, 11975072, 11875102, and 11835009), the China Manned Space Program (Grant No. CMS-CSST-2025-A02), and the National 111 Project (Grant No. B16009). Xiao Guo acknowledges the fellowship of China National Postdoctoral Program for Innovative Talents (Grant No. BX20230104).